### v1

# Adaptive Quantum Control

Abstract
This paper presents a model-free, learning-based adaptive controller for the density tracking problem in a two-level Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) quantum system. The adaptive controller is based on the continuous-time retrospective cost adaptive control. To preserve the geometric properties of the quantum system, an adaptive PID controller driven and optimized by Ulhmann’s fidelity is used. The proposed controller is validated in simulation for a low and a high-entropy density-tracking problem.
keywords: adaptive control, quantum control, density-tracking.
I INTRODUCTION
Quantum dynamical systems have several applications in sensing, metrology, communication, computation, and the internet [giovannetti2011advances, toth2014quantum, taylor2016quantum, cacciapuoti2020entanglement, nielsen2001quantum]. The reliable operation of quantum circuits, such as quantum logic gates, requires the precise and predictable behavior of quantum dynamical systems, much like how classical digital circuits depend on the accurate design of their fundamental electrical components, such as resistors and transistors [barenco1995elementary]. In quantum computing, the basic unit of information is the qubit, which, unlike classical bits limited to two distinct states (00 or $1$), can exist in a superposition of both states, enabling quantum algorithms to perform computations exponentially faster than classical systems [grover1996fast, shor1999polynomial]. The successful realization of quantum devices depends on the precise control of quantum dynamical systems.
Quantum control involves manipulating quantum systems to achieve specific outcomes. It has applications in fields such as quantum computing, communication, and sensing. This is achieved through external controls, such as electromagnetic fields, to direct the evolution of quantum systems. The nature of the desired quantum event in a control problem depends on the quantum problem objective. For instance, in applications such as quantum computing, metrology, communication, and optical systems [yamamoto2024engineering, zaletel2021preparation], a low entropy state is desired to reduce the randomness and uncertainties from the output, whereas high entropy is desired when high randomness is needed in quantum phenomena. Applications such as heat spin engines, quantum cryptography, and quantum thermodynamics simulation [von2019spin, Sehgal2021, somhorst2023quantum], can be conceived as quantum control problems of this category. Quantum control strategies often rely on optimizing the system’s time evolution, governed by quantum operators such as Hamiltonians and density matrices. The Schrödinger equation is used for closed systems, while open systems interacting with the environment follow the LGKS equation [brif2010control]. Recent advancements have improved control techniques for both isolated and open quantum systems [dong2010quantum, cong2014control]. However, due to fundamental physical limitations of quantum state measurement and uncertainty in the knowledge of quantum dynamics, the performance achievable by model-based control systems is limited.
Adaptive quantum control addresses the challenges posed by fundamental physical limitations in quantum state measurement and uncertainties in quantum dynamics. This control strategy adjusts its parameters based on feedback from the quantum system, compensating for variations in system parameters and external disturbances[Kosut2003, kosut2013adaptive, Liu2020]. By incorporating online learning algorithms, adaptive control reduces the impact of incomplete system knowledge and measurement limitations [egger2014adaptive] This approach is significant for practical quantum technologies, including quantum error correction and quantum information processing, where uncertainties affect performance.
Despite these advancements, existing adaptive controllers primarily address parameter uncertainty while still requiring some knowledge of the quantum system model, which can be difficult to obtain in complex quantum systems. This fundamental assumption enables adaptive controllers in the literature to focus mainly on managing parameter uncertainties [egger2014adaptive, Koswara2014, Hu2016].
This paper considers the problem of designing an adaptive controller for an open quantum system without any prior knowledge of the system’s dynamics. In this preliminary work, we consider Uhlmann’s fidelity as the error metrics [UHLMANN1976273]. Uhlmann’s fidelity is a unique real scalar metric that allows one to compare the similarity of two quantum states. In this paper, we design a continuous-time adaptive controller based on the Retrospective Cost Adaptive Control (RCAC) Technique [ali2015retrospective, Goel2015, Sobolic2016, poudel2023learning] for the tracking problem of an open quantum system whose dynamics are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) equation [manzano2020short]. The contributions of this paper are thus 1) the design of a continuous-time adaptive controller optimized by RCAC for the open quantum system described by the LGKS equation without prior knowledge of the system’s dynamics, 2) the design of an adaptive PID controller driven by the Uhlmann’s fidelity, and 3) numerical validation of the proposed controller in a low and high entropy state-tracking problem.
The paper is organized as follows. Section II describes the LGKS equation in the context of a controlled quantum open system. Section III presents the mathematical foundations for continuous time RCAC with forgetting factors. Section IV describes the numerical simulation that validate the application of the adaptive controller in a two-level open system. Finally, the paper concludes with a discussion of results and future research directions in section V.
II Problem Formulation
Consider the Lindblad-Gorini-Kossakowski-Sudarshan equation [manzano2020short]
$\dot{\rho}$ $=-i\hbar[H,\rho]+\sum_{i=1}^{n}\left(L_{i}\rho L_{i}^{\rm H}-\dfrac{1}{2}\{L_{i}^{\rm H}L_{i},\rho\}\right),$ (1)
where $\rho\in\mathbb{C}^{2\times 2}$ is the density matrix, the system Hamiltonian $H\in\mathbb{C}^{2\times 2}$ is Hermitian, $n$ is the number of damping terms, and for $i\in(1,2,\ldots,n),$ the matrices $L_{i}\in{\mathbb{C}}^{2\times 2}$ are the jump operators. Without loss of generality, the units are chosen such that $\hbar=1.$ The diagonal elements of $\rho$ represent the probability of a quantum event. Thus, in the case where $\rho\in{\mathbb{C}}^{2\times 2},$ (1) is called a two-level quantum dynamic system.
It is shown in [breuer2002theory, p 123] that, if ${\rm tr}\,\rho(0)=1,$ then, for all $t>0,$ $\rho(t)$ given by (1) satisfies ${\rm tr}\,(\rho(t))=1.$ Furthermore, for all $t\geq 0,$ $\rho(t)$ is positive semidefinite and Hermitian.
As shown in [altafini2012modeling], the Hamilotonian $H$ can be decomposed as
$H$ $=H_{0}+H_{1}u,$ (2)
where $H_{0}\in{\mathbb{C}}^{2\times 2}$ is the free Hamiltonian, $H_{1}\in{\mathbb{C}}^{2\times 2}$ is the control Hamiltonian, and $u$ is the control signal to be designed. In this paper, we consider the case where $u\in{\mathbb{R}},$ which ensures that the Hamiltonian $H$ is Hermitian at all time instants.
The control objective is to design an adaptive feedback control law such that the density matrix $\rho$ converges to the desired density matrix $\rho_{\rm d}.$
II-A Density Error
As shown in [jones2024controlling], the Uhlmann–Jozsa fidelity between two density matrices $\rho_{1},\rho_{2}$ is defined as
$F(\rho_{1},\rho_{2})$ $\stackrel{{\scriptstyle\triangle}}{{=}}\left({\rm tr}\,\left(\sqrt{\sqrt{\rho_{1}}\,\rho_{2}\,\sqrt{\rho_{1}}}\right)\right)^{2}.$ (3)
It follows from Theorem A.5 that $F(\rho_{1},\rho_{2})\in[0,1].$ Furthermore, $F(\rho_{1},\rho_{2})=1$ if and only if $\rho_{1}=\rho_{2}.$ Finally, as shown in [jozsa1994fidelity], in the case where $\rho_{1},\rho_{2}\in\mathbb{C}^{2\times 2},$ the Uhlmann–Jozsa fidelity is given by
$F(\rho_{1},\rho_{2})$ $={\rm tr}\,(\rho_{1}\rho_{2})+2\sqrt{\det(\rho_{1})\det(\rho_{2})}$ (4)
In this paper, we consider the density error
$e$ $\stackrel{{\scriptstyle\triangle}}{{=}}1-F(\rho,\rho_{\rm{d}})\in[0,1],$ (5)
where $\rho_{\rm{d}}\in\mathbb{C}^{2\times 2}$ is the desired density
II-B Adaptive PID Control
An adaptive proportional-integral-derivative controller can be written as
$u(t)$ $=k_{\rm p}(t)e(t)+k_{\rm i}(t)\int_{0}^{t}e(\tau){\rm d}\tau+k_{\rm d}(t)\dot{e}(t),$ (6)
where $k_{\rm p}(t),k_{\rm i}(t),$ and $k_{\rm d}(t)$ are the adaptive proportional, integral, and derivative gains. Note that the controller can be written as
$u(t)=\Phi(t)\theta(t),$ (7)
where
$\Phi(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}e(t)&\gamma(t)&\dot{e}(t)\end{bmatrix},\quad\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}k_{\rm p}(t)\\ k_{\rm i}(t)\\ k_{\rm d}(t)\end{bmatrix}.$ (8)
Note that integral state $\gamma(t)$ satisfies $\dot{\gamma}(t)=e(t).$
The objective of the adaptive controller is to update the controller gains and thus compute the control signal $u(t)$ such that the density error $e(t)$ converges to zero. In this paper, we use the continuous-time retrospective cost adaptive control algorithm, described in the next section, to update the controller gains.
III Continuous Time Retrospective Cost Adaptive Control
Consider a dynamic system
$\dot{x}(t)$ $=f(x(t),u(t)),$ (9)
$y(t)$ $=g(x(t)),$ (10)
where $x(t)\in\mathbb{C}^{l_{x}}$ is the state, $u(t)\in\mathbb{C}^{l_{u}}$ is the input, $y(t)\in\mathbb{C}^{l_{y}}$ is the measured output, and the vector functions $f$ and $g$ are the dynamics and the output maps. Define the performance variable
$z(t)\stackrel{{\scriptstyle\triangle}}{{=}}y(t)-r(t),$ (11)
where $r(t)$ is the exogenous reference signal. The objective is to design an adaptive output feedback control law to ensure that $z(t)\to 0.$
Consider a linearly parameterized control law
$u(t)=\Phi(t)\theta(t),$ (12)
where the regressor matrix $\Phi(t)\in\mathbb{C}^{l_{u}\times l_{\theta}}$ contains the measured data and the vector $\theta(t)\in\mathbb{C}^{l_{\theta}}$ contains the controller gains to be optimized. Various linear parameterizations of MIMO controllers are described in [goel_2020_sparse_para].
Next, using (11), define the retrospective performance
$\hat{z}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}z(t)+\Phi_{\rm f}(t)\hat{\theta}(t)-u_{\rm f}(t),$ (13)
where $\hat{\theta}(t)$ is the controller gain to be optimized and the filtered regressor $\Phi_{\rm f}(t)$ and the filtered control $u_{\rm f}(t)$ are defined as
$\Phi_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[\Phi(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (14)
$u_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[u(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (15)
where $G_{\rm f}(s)$ is a dynamic filter.
The filtered signals $\Phi_{\rm f}(t)$ and $u_{\rm f}(t)$ are computed in the time domain, as shown below. Let $(A_{\rm f},B_{\rm f},C_{\rm f},D_{\rm f})$ be a realization of $G_{\rm f}(s).$ Then,
$\dot{x}_{\Phi}$ $=A_{\rm f}x_{\phi}+B_{\rm f}\Phi,$ (16)
$\Phi_{\rm f}$ $=C_{\rm f}x_{\phi}+D_{\rm f}\Phi,$ (17)
and
$\dot{x}_{u}$ $=A_{\rm f}x_{u}+B_{\rm f}u,$ (18)
$u_{\rm f}$ $=C_{\rm f}x_{u}+D_{\rm f}u.$ (19)
Next, define the retrospective cost
$J(t,\hat{\theta})$ $=\int_{0}^{t}\Big{(}e^{-\lambda(t-\tau)}\hat{z}(\tau)^{\rm T}R_{z}\hat{z}(\tau)$
$\quad+(\Phi(\tau)\hat{\theta})^{\rm T}R_{u}(\Phi(\tau)\hat{\theta})\Big{)}{\rm d}\tau$
$\quad+e^{-\lambda t}\hat{\theta}^{T}R_{\theta}\hat{\theta},$ (20)
where $R_{z},R_{u},$ and $R_{\theta}$ are positive definite weighting matrices of appropriate dimensions, and $\lambda<0$ is the exponential forgetting factor.
Proposition III.1.
Consider the cost function $J(t,\hat{\theta})$ given by (20). For all $t\geq 0,$ define the minimizer of $J(t,\hat{\theta})$ by
$\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\underset{\hat{\theta}\in{\mathbb{R}}^{l_{\theta}}}{\operatorname{argmin}}J(t,\hat{\theta}).$ (21)
Then, for all $t\geq 0,$ the minimizer satisfies
$\dot{\theta}$ $=-P\Phi_{\rm f}^{\rm T}R_{z}(z+\Phi_{\rm f}\theta-u_{\rm f})-P\Phi^{\rm T}R_{u}\Phi\theta,$ (22)
$\dot{P}$ $=\lambda P-P\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right)P$ (23)
where $P(0)=R_{\theta}^{-1}$ and $\theta(0)=0.$
Proof.
Note that the cost function (20) can be written as
$J(t,\hat{\theta})$ $=\hat{\theta}^{T}A(t)\hat{\theta}+2\hat{\theta}^{T}b(t)+c(t),$ (24)
where
$A(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\int_{0}^{t}e^{-\lambda(t-\tau)}\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right){\rm d}\tau+e^{-\lambda t}R_{\theta},$
$b(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\int_{0}^{t}e^{-\lambda(t-\tau)}\Phi_{\rm f}^{\rm T}R_{z}(z-u_{\rm f}){\rm d}\tau,$
$c(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\int_{0}^{t}e^{-\lambda(t-\tau)}(z-u_{\rm f})^{\rm T}R_{z}(z-u_{\rm f}){\rm d}\tau.$
The matrices $A(t)$ and $b(t)$ satisfy
$\dot{A}$ $=-\lambda A+\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi,$
$\dot{b}$ $=-\lambda b+\Phi_{\rm f}^{\rm T}R_{z}(z-u_{\rm f}),$
where $A(0)=R_{\theta}$ and $b(0)=0.$
Next, define, for all $\geq 0,$ $P(t)\stackrel{{\scriptstyle\triangle}}{{=}}A(t)^{-1}.$ Using the fact that $P(t)A(t)=I,$ it follows that
$\dot{P}(t)=-P(t)\dot{A}(t)P(t),$
and thus
$\dot{P}=\lambda P-P\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right)P.$
Finally, note that the minimizer of (24) is given by
$\theta(t)=-A(t)^{-1}b(t)=-P(t)b(t),$
and thus
$\dot{\theta}(t)$ $=-\dot{P}b-P\dot{b}$
$=-P\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right)PP^{-1}\theta$
$\quad-P\Phi_{\rm f}^{\rm T}R_{z}(z-u_{\rm f})$
$=-P\Phi_{\rm f}^{\rm T}R_{z}(z+\Phi_{\rm f}\theta-u_{\rm f})-P\Phi^{\rm T}R_{u}\Phi\theta.$
The control law is thus
$u(t)=\Phi(t)\theta(t),$ (25)
where $\theta(t)$ is given by (22).
IV Numerical Simulation
In this paper, we consider a two-level system, with dynamics given by the LGKS (1), with a single damped term, that is, $n=1$. The control system architecture is shown in Figure 1.
$F(\rho,\rho_{\rm{d}})$$\begin{array}[]{c}\text{Adaptive}\\ \text{Controller}\end{array}$ $\mathcal{L}\mathcal{G}\mathcal{K}\mathcal{S}$$\rho_{\rm{d}}$$e$$u$$\rho$
Figure 1: Adaptive control architecture to track desired density matrices in the LGKS system.
To simulate the LKGS equation, we set
$H_{0}=0.5\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},H_{1}=0.5\begin{bmatrix}0&1\\ 1&0\end{bmatrix},L_{1}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix},$ (26)
where $H_{0}$ and $H_{1}$ are Pauli matrices that, as described in [d2001optimal, kamakari2022digital], satisfy the controllability conditions in the problem of controlling a two-level quantum system, and $L_{1}$ is selected as a non-Hermitian matrix, as suggested in [verstraete2009quantum, schlimgen2022quantum].
The initial density matrix $\rho(0)$ is assumed to be
$\rho_{0}$ $=\begin{bmatrix}0.4&0.1+0.3\imath\\ 0.1-0.3\imath&0.6\end{bmatrix}.$ (27)
Note that $\rho(0)$ is randomly generated such that ${\rm tr}\,\rho(0)=1$ and it is Hermitian. As shown in [moses2015creation] and [xu2021multicriticality], we consider the problem of moving the quantum state $\rho(0)$ to a low and a high desired entropy states $\rho_{\rm d},$ where the Von Neumann entropy of a quantum state $\rho$ is given by
$S(\rho)\stackrel{{\scriptstyle\triangle}}{{=}}-{\rm tr}\,(\rho\,\rm{ln}(\rho)).$ (28)
First, we consider the problem of driving the quantum state $\rho(t)$ to
$\rho_{\rm d}=\begin{bmatrix}0.8571&0.2857+0.1429\imath\\ 0.2857-0.1429\imath&0.1429\end{bmatrix}.$ (29)
Note that $S(\rho_{\rm d})=0.1013,$ and thus, this desired state is a low-entropy state. Furthermore, note that $\rho_{\rm d},$ given by (29) is an equilibrium point of (1) in the case where $u(t)\equiv 1.$
With the desired density $\rho_{\rm d}$ given by (29) as the reference for the controller, we first tune the RCAC hyperparameters. In particular, in RCAC, we set $R_{z}=1,$ $R_{u}=1,$ and $\lambda=0.01$; and optimize $P_{0}$ and the filter coefficient $\beta,$ where $G_{\rm f}(s)=\dfrac{1}{s+\beta},$ as described below. We optimize only $P_{0}$ and the filter coefficient since it has been observed in several numerical experiments that the choice of $P_{0}$ and the filter coefficient has the most dominant impact on the closed-loop performance. For each pair of $(P_{0},\beta)\in\{(10^{-5},10^{-4},\ldots,10^{10})\times(0,1,2,5,100,2000)\},$ we apply the RCAC algorithm to update the adaptive PID controller to drive the quantum state from $\rho_{0}$, given by (27) to $\rho_{\rm d}$, given by (29). Each simulation is run for 200 seconds. At the end of each simulation, the hyperparameter cost
$J_{h}\stackrel{{\scriptstyle\triangle}}{{=}}\int_{190}^{200}|z(\tau)|{\rm d}\tau$ (30)
is computed. Figure 2 shows the contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The tuning pair with the minimum value of $J_{h}$ is selected for all the numerical simulations in the rest of the paper. In particular, in all the examples, we set $P_{0}=10^{-3}\times I_{3},\beta=2000.$
Figure 2: Contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The pair $(P0,\beta)$ that minimizes $J_{h}$ is used in all numerical simulations.
IV-A Low-entropy Density
With the selected RCAC hyperparameters, Figure 3 shows the absolute value of the density error $e(t)$ between the system state and the desired low-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller. Figure 4 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state. Figure 5 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere. Note that the quantum state $\rho$ satisfies $x=2\,\rm{Real}(\rho_{1,2}),$ $y=2\,\rm{Imag}(\rho_{1,2}),$ and $z=\rho_{1,1}-\rho_{2,2}$ on the Bloch sphere [wie2020two].
Figure 3: Low-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
Figure 4: Low-entropy density tracking. The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 5: Low-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
IV-B High-entropy Density
Next, we consider the problem of driving the quantum state $\rho(t)$ to
$\rho_{\rm d}=\begin{bmatrix}0.5168&0.0971+0.0460i\\ 0.0971-0.0460i&0.4832\end{bmatrix}.$ (31)
Note that $S(\rho_{\rm d})=0.6693,$ and thus, this desired state is a high-entropy state. Furthermore, note that $\rho_{\rm d},$ given by (31) is an equilibrium point of (1) in the case where $u(t)\equiv 10.$
Figure 6 shows the absolute value of the density error $e(t)$ between the system state and the desired high-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller. Figure 7 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state. Figure 8 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
The two examples described above suggest that the adaptive PID controller updated by the RCAC algorithm can effectively track an arbitrary desired density matrix $\rho_{\rm d}$ without requiring any model of the underlying dynamics of the density matrix.
Figure 6: High-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
Figure 7: High-entropy density tracking. The first two subplot shows the real and imaginary part of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 8: High-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
V Conclusions
This paper presented an online learning-based, model-free adaptive controller for tracking a desired density matrix in a two-level Lindblad-Gorini-Kossakowski-Sudarshan quantum system. In particular, the adaptive controller uses the retrospective cost adaptive control algorithm. Since the classical RCAC was developed for a discrete-time system, the paper presented a continuous-time extension of RCAC. An adaptive PID controller structure was chosen to generate a control signal that preserves the LKGS solution’s geometric structure. Uhlmann-Jozsa fidelity was used as an error metric to drive the adaptive PID controller. Numerical simulations show that the proposed controller effectively tracks desired low and high-entropy states without requiring a model of the underlying system dynamics.
Future work will extend the technique to the case where the input is a Hermitian matrix and the density coherence is used as an error metric. This will allow higher-dimensional parameterization of the adaptive PID controller, which may yield better tracking performance.
[…]
Appendix
A Some Useful Facts
This appendix reviews some useful definitions and facts about Hermitian matrices used in the LKGS system.
Definition A.1.
Let $A\in{\mathbb{C}}^{n\times n}.$ $A$ is a Hermitian matrix if $A=\overline{A^{\rm T}},$ where $\overline{A^{\rm T}}$ is the conjugate transpose of $A.$
The conjguate transpose of $A$ is denoted by $A^{\rm H}.$ If $A$ is Hermitian, then $A=A^{\rm H}.$
Definition A.2.
Let $A,B\in{\mathbb{C}}^{n\times n}.$ Then, the commutator $[A,B]$ is defined as
$[A,B]\stackrel{{\scriptstyle\triangle}}{{=}}AB-BA,$ (32)
and the anticommutator $\{A,B\}$ is defined as
$\{A,B\}\stackrel{{\scriptstyle\triangle}}{{=}}AB+BA.$ (33)
Fact A.1.
Let $A,B\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $-\imath[A,B]$ is Hermitian.
Proof.
Note that $(-i[A,B])^{\rm H}=(-\imath)^{\rm H}[A,B]^{\rm H}=\imath\left(AB-BA\right)^{\rm H}=\imath(A^{\rm H}B^{\rm H}-B^{\rm H}A^{\rm H})=\imath(BA-AB)=-\imath(AB-BA)=-\imath[A,B]$. Thus, $-\imath[A,B]$ is Hermitian. ∎
Fact A.2.
Let $U\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $U^{\rm H}U$ and $UU^{\rm H}$ are Hermitian.
Proof.
Note that $(U^{\rm H}U)^{\rm H}=U^{\rm H}(U^{\rm H})^{\rm H}=U^{\rm H}U,$ which implies that $UU^{\rm H}$ is Hermitian. A similar argument proves that $U^{\rm H}U$ is also Hermitian. ∎
Fact A.3.
Let $A,U\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $UAU^{\rm H}$ is Hermitian.
Proof.
Note that $(UAU^{\rm H})^{\rm H}=(U^{\rm H})^{\rm H}(UA)^{\rm H}=UA^{\rm H}U^{\rm H}=UAU^{\rm H},$ which implies that $UAU^{\rm H}$ is Hermitian. ∎
Fact A.4.
Let $A,B,C\in{\mathbb{C}}^{n\times n}$ be Hermitian. Suppose $B$ satisfies
$\dot{B}=-\imath[A,B]+\sum_{i=1}^{N}\left(CBC^{\rm H}-\dfrac{1}{2}\{C^{\rm H}C,B\}\right).$ (34)
Then, for all $t\geq 0,$ ${\rm tr}\,B$ is a constant.
Proof.
Note that $\dfrac{{\rm d}}{{\rm d}t}{\rm tr}\,(B)={\rm tr}\,(\dot{B})$. Next, note that ${\rm tr}\,(-\imath[A,B])=-\imath{\rm tr}\,(AB-BA)=-\imath{\rm tr}\,(AB)+\imath{\rm tr}\,(BA)=-\imath{\rm tr}\,(AB)+\imath{\rm tr}\,(AB)=0.$ Using the cyclic property of the trace, it follows that
${\rm tr}\,\left(\sum_{i}^{N}\left(CBC^{\rm H}-\dfrac{1}{2}(C^{\rm H}CB+BC^{\rm H}C)\right)\right)$
$=\sum_{i}^{N}{\rm tr}\,(CBC^{\rm H})-\dfrac{1}{2}\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)-\dfrac{1}{2}\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)$
$=\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)-\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)=0,$
which completes the proof. ∎
Fact A.5.
Let $\rho\in\mathbb{C}^{n\times n}$ and $\sigma\in\mathbb{C}^{n\times n}$ be two Hermitian PSD matrices with traces equal to $1$. Then,
$1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$
Proof.
Since $\rho$ and $\sigma$ are PSD Hermitian. Then, $\rho^{{\rm H}}=\rho$, $\sigma^{{\rm H}}=\sigma$, $x^{{\rm H}}\rho x\geq 0$ and $x^{{\rm H}}\sigma x\geq 0$ for all $x\in\mathbb{C}.$ Furthermore, the unique square root of a PSD Hermitian matrix is also Hermitian and PSD [horn2012matrix, p 440]. This implies that, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\left(\sqrt{(}\rho)\sigma\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\sigma^{{\rm H}}\sqrt{\rho}^{{\rm H}}=\sqrt{\rho}\sigma\sqrt{\rho}.$ Define, $y\stackrel{{\scriptstyle\triangle}}{{=}}\sqrt{\rho}x\in\mathbb{C}^{n}.$ Then, $x^{{\rm H}}\sqrt{\rho}\sigma\sqrt{\rho}x=y^{{\rm H}}\sigma y\geq 0,$ which implies that $\sqrt{\rho}\sigma\sqrt{\rho}\geq 0.$ Therefore, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)$ is Hermitian, and $\sqrt{\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)}$ is Hermitian and unique. From the Cauchy-Schwarz inequality and the cyclic property for traces, $|{\rm tr}\,(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}})|^{2}=|{\rm tr}\,(\sqrt{\sqrt{\rho}\sqrt{\rho}\sigma})|^{2}=|{\rm tr}\,(\sqrt{\rho\sigma})|^{2}\leq{\rm tr}\,(\rho){\rm tr}\,(\sigma)\leq 1$. Thus, $1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\rho_{\rm{d}}\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$

### v2

@@
-Adaptive Quantum Control
+Adaptive Fidelity-Based Density Tracking for Open Quantum Systems
 Abstract
-This paper presents a model-free, learning-based adaptive controller for the density tracking problem in a two-level Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) quantum system.
-The adaptive controller is based on the continuous-time retrospective cost adaptive control.
-To preserve the geometric properties of the quantum system, an adaptive PID controller driven and optimized by Ulhmann’s fidelity is used.
-The proposed controller is validated in simulation for a low and a high-entropy density-tracking problem.
-keywords: adaptive control, quantum control, density-tracking.
+This paper presents an online learning-based adaptive control framework for density-matrix tracking in a two-level Lindblad–Gorini–Kossakowski–Sudarshan (LGKS) quantum system, in which the feedback control law does not require prior knowledge of the system Hamiltonian or dissipative operators.
+The adaptive controller is based on a continuous-time formulation of retrospective cost adaptive control (RCAC).
+To preserve the geometric structure of the quantum-state evolution, an adaptive PID controller driven by Uhlmann’s fidelity is employed.
+The proposed approach is validated in numerical simulations for both low-entropy and high-entropy density-tracking tasks, and robustness to measurement noise in the feedback path is investigated.
+keywords: Quantum feedback control, open quantum systems, online learning, adaptive PID control, Lindblad dynamics
 I INTRODUCTION
-Quantum dynamical systems have several applications in sensing, metrology, communication, computation, and the internet [giovannetti2011advances, toth2014quantum, taylor2016quantum, cacciapuoti2020entanglement, nielsen2001quantum].
-The reliable operation of quantum circuits, such as quantum logic gates, requires the precise and predictable behavior of quantum dynamical systems, much like how classical digital circuits depend on the accurate design of their fundamental electrical components, such as resistors and transistors [barenco1995elementary].
-In quantum computing, the basic unit of information is the qubit, which, unlike classical bits limited to two distinct states (00 or $1$), can exist in a superposition of both states, enabling quantum algorithms to perform computations exponentially faster than classical systems [grover1996fast, shor1999polynomial].
-The successful realization of quantum devices depends on the precise control of quantum dynamical systems.
-Quantum control involves manipulating quantum systems to achieve specific outcomes.
-It has applications in fields such as quantum computing, communication, and sensing.
-This is achieved through external controls, such as electromagnetic fields, to direct the evolution of quantum systems.
-The nature of the desired quantum event in a control problem depends on the quantum problem objective.
-For instance, in applications such as quantum computing, metrology, communication, and optical systems [yamamoto2024engineering, zaletel2021preparation], a low entropy state is desired to reduce the randomness and uncertainties from the output, whereas high entropy is desired when high randomness is needed in quantum phenomena.
-Applications such as heat spin engines, quantum cryptography, and quantum thermodynamics simulation [von2019spin, Sehgal2021, somhorst2023quantum], can be conceived as quantum control problems of this category.
-Quantum control strategies often rely on optimizing the system’s time evolution, governed by quantum operators such as Hamiltonians and density matrices.
-The Schrödinger equation is used for closed systems, while open systems interacting with the environment follow the LGKS equation [brif2010control].
-Recent advancements have improved control techniques for both isolated and open quantum systems [dong2010quantum, cong2014control].
-However, due to fundamental physical limitations of quantum state measurement and uncertainty in the knowledge of quantum dynamics, the performance achievable by model-based control systems is limited.
-Adaptive quantum control addresses the challenges posed by fundamental physical limitations in quantum state measurement and uncertainties in quantum dynamics.
-This control strategy adjusts its parameters based on feedback from the quantum system, compensating for variations in system parameters and external disturbances[Kosut2003, kosut2013adaptive, Liu2020].
-By incorporating online learning algorithms, adaptive control reduces the impact of incomplete system knowledge and measurement limitations [egger2014adaptive] This approach is significant for practical quantum technologies, including quantum error correction and quantum information processing, where uncertainties affect performance.
-Despite these advancements, existing adaptive controllers primarily address parameter uncertainty while still requiring some knowledge of the quantum system model, which can be difficult to obtain in complex quantum systems.
-This fundamental assumption enables adaptive controllers in the literature to focus mainly on managing parameter uncertainties [egger2014adaptive, Koswara2014, Hu2016].
-This paper considers the problem of designing an adaptive controller for an open quantum system without any prior knowledge of the system’s dynamics.
-In this preliminary work, we consider Uhlmann’s fidelity as the error metrics [UHLMANN1976273].
-Uhlmann’s fidelity is a unique real scalar metric that allows one to compare the similarity of two quantum states.
-In this paper, we design a continuous-time adaptive controller based on the Retrospective Cost Adaptive Control (RCAC) Technique [ali2015retrospective, Goel2015, Sobolic2016, poudel2023learning] for the tracking problem of an open quantum system whose dynamics are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) equation [manzano2020short].
-The contributions of this paper are thus 1) the design of a continuous-time adaptive controller optimized by RCAC for the open quantum system described by the LGKS equation without prior knowledge of the system’s dynamics, 2) the design of an adaptive PID controller driven by the Uhlmann’s fidelity, and 3) numerical validation of the proposed controller in a low and high entropy state-tracking problem.
-The paper is organized as follows.
-Section II describes the LGKS equation in the context of a controlled quantum open system.
-Section III presents the mathematical foundations for continuous time RCAC with forgetting factors.
-Section IV describes the numerical simulation that validate the application of the adaptive controller in a two-level open system.
-Finally, the paper concludes with a discussion of results and future research directions in section V.
+Quantum dynamical systems have several applications in sensing, metrology, communication, computation, and the internet [10, 30, 29, 6, 23].
+The reliable operation of quantum circuits, such as quantum logic gates, requires the precise and predictable behavior of quantum dynamical systems, much like how classical digital circuits depend on the accurate design of their fundamental electrical components, such as resistors and transistors [3].
+In quantum computing, the basic unit of information is the qubit, which, unlike classical bits limited to two distinct states (0 or $1$), can exist in a superposition of both states, enabling quantum algorithms to perform computations exponentially faster than classical systems [12, 26].
+The successful realization of quantum devices thus depends on the precise control of quantum dynamical systems.
+Quantum control concerns the manipulation of quantum systems to achieve specific dynamical or state-transfer objectives and plays a central role in quantum computing, communication, and sensing.
+Such control is typically realized through externally applied fields, for example, electromagnetic control inputs, that steer the system’s evolution.
+The desired quantum state depends on the application.
+In quantum computing, metrology, communication, and optical systems [35, 36], low-entropy (high-purity) states are often desirable to preserve coherence and reduce uncertainty, whereas in other settings, high-entropy or mixed states may be of interest when intrinsic randomness or thermal effects are central to the application.
+Examples of the latter include heat engines based on spin systems, quantum cryptography, and quantum thermodynamics simulations [32, 25, 28].
+Quantum control strategies typically seek to shape the system’s time evolution, governed by quantum operators such as Hamiltonians and density matrices.
+Closed quantum systems evolve according to the Schrödinger equation, whereas open quantum systems interacting with an environment are commonly modeled using the Lindblad–Gorini–Kossakowski–Sudarshan (LGKS) equation [5].
+Recent years have seen significant progress in control techniques for both isolated and open quantum systems [8, 7].
+In practice, the performance of model-based quantum control strategies is limited by uncertainty in the underlying system dynamics, unmodeled environmental effects, and constraints on the accuracy and availability of quantum state measurements.
+These challenges motivate the development of adaptive quantum control methods that adjust controller parameters online based on system feedback to compensate for modeling errors, parameter drift, and disturbances [17, 18, 20].
+By incorporating online learning mechanisms, adaptive control can mitigate the impact of incomplete system knowledge and imperfect measurements [9].
+Such approaches are relevant for practical quantum technologies, including quantum error correction and quantum information processing, where uncertainties and environmental interactions directly affect performance.
+Despite these advances, most existing adaptive quantum control approaches primarily address parametric uncertainty while still requiring partial knowledge of the system model, such as the Hamiltonian structure or dissipative operators, which may be difficult to obtain accurately for complex quantum systems.
+Consequently, much of the existing literature focuses on robustness to parameter variations rather than control design in the absence of an explicit dynamical model [9, 19, 14].
+This motivates the investigation of control architectures in which the feedback law itself does not rely on detailed knowledge of the system dynamics.
+This paper considers the problem of designing an adaptive controller for an open quantum system in which the control law does not require prior knowledge of the system Hamiltonian or dissipative operators.
+Specifically, we develop a continuous-time adaptive controller based on the Retrospective Cost Adaptive Control (RCAC) framework [1, 11, 27, 24] for density-matrix tracking in open quantum systems governed by the Lindblad–Gorini–Kossakowski–Sudarshan (LGKS) equation [21].
+In this preliminary study, Uhlmann’s fidelity is adopted as the scalar performance metric driving the adaptation [31], providing a geometrically meaningful measure of similarity between quantum states.
+The main contributions of this paper are:
+  • the formulation of a continuous-time RCAC-based adaptive controller for LGKS quantum dynamics without requiring a plant model in the control law,
+  • the design of an adaptive PID feedback structure driven by fidelity-based error, and
+  • numerical validation of the proposed approach for tracking both low-entropy and high-entropy target states in a two-level open quantum system.
+The remainder of the paper is organized as follows.
+Section II describes the LGKS equation in the context of a controlled open quantum system.
+Section III presents the mathematical foundations of continuous-time RCAC with forgetting factors.
+Section IV presents numerical simulations validating the proposed adaptive controller on a two-level quantum system.
+Finally, Section V concludes the paper and outlines directions for future work.
 II Problem Formulation
-Consider the Lindblad-Gorini-Kossakowski-Sudarshan equation [manzano2020short]
-$\dot{\rho}$ $=-i\hbar[H,\rho]+\sum_{i=1}^{n}\left(L_{i}\rho L_{i}^{\rm H}-\dfrac{1}{2}\{L_{i}^{\rm H}L_{i},\rho\}\right),$ (1)
-where $\rho\in\mathbb{C}^{2\times 2}$ is the density matrix, the system Hamiltonian $H\in\mathbb{C}^{2\times 2}$ is Hermitian, $n$ is the number of damping terms, and for $i\in(1,2,\ldots,n),$ the matrices $L_{i}\in{\mathbb{C}}^{2\times 2}$ are the jump operators.
+We consider the Lindblad-Gorini-Kossakowski-Sudarshan system as described in [21], which is governed by
+${\dot{\rho}}$ $=-i\hbar[H,{{\rho}}]+\sum_{i=1}^{m}\left(L_{i}{{\rho}}L_{i}^{\rm H}-\dfrac{1}{2}\{L_{i}^{\rm H}L_{i},{{\rho}}\}\right),$ (1)
+where ${{\rho}}\in\mathbb{C}^{n\times n}$ is the density matrix, the system Hamiltonian $H\in\mathbb{C}^{n\times n}$ is Hermitian, $m$ is the number of damping terms, and for $i\in(1,2,\ldots,m),$ the matrices $L_{i}\in{\mathbb{C}}^{n\times n}$ are the jump operators.
 Without loss of generality, the units are chosen such that $\hbar=1.$ The diagonal elements of $\rho$}} represent the probability of a quantum event.
-Thus, in the case where $\rho\in{\mathbb{C}}^{2\times 2},$ (1) is called a two-level quantum dynamic system.
-It is shown in [breuer2002theory, p 123] that, if ${\rm tr}\,\rho(0)=1,$ then, for all $t>0,$ $\rho(t)$ given by (1) satisfies ${\rm tr}\,(\rho(t))=1.$ Furthermore, for all $t\geq 0,$ $\rho(t)$ is positive semidefinite and Hermitian.
-As shown in [altafini2012modeling], the Hamilotonian $H$ can be decomposed as
+It is shown in [4, p 123] that, if ${\rm tr}\,{{\rho}}(0)=1,$ then, for all $t>0,$ ${{\rho}}(t)$ given by (1) satisfies ${\rm tr}\,({{\rho}}(t))=1.$ Furthermore, for all $t\geq 0,$ ${{\rho}}(t)$ is positive semidefinite and Hermitian.
+In this paper, we consider a two-level quantum dynamic system, that is, $n=2$ and thus ${{\rho}}\in{\mathbb{C}}^{2\times 2}.$ As shown in [2], the Hamilotonian $H$ can be decomposed as
 $H$ $=H_{0}+H_{1}u,$ (2)
-where $H_{0}\in{\mathbb{C}}^{2\times 2}$ is the free Hamiltonian, $H_{1}\in{\mathbb{C}}^{2\times 2}$ is the control Hamiltonian, and $u$ is the control signal to be designed.
-In this paper, we consider the case where $u\in{\mathbb{R}},$ which ensures that the Hamiltonian $H$ is Hermitian at all time instants.
-The control objective is to design an adaptive feedback control law such that the density matrix $\rho$ converges to the desired density matrix $\rho_{\rm d}.$
+where $H_{0}\in\mathbb{C}^{2\times 2}$ is the free Hamiltonian, $H_{1}\in\mathbb{C}^{2\times 2}$ is the control Hamiltonian, and $u\in\mathbb{R}$ is the scalar control input.
+The restriction $u\in\mathbb{R}$ ensures that $H$ remains Hermitian at all times.
+As shown in Figure 1, the objective is to design an adaptive feedback control law that drives the density matrix $\rho$ to a desired target state $\rho_{\rm d}$ without requiring prior knowledge of $H_{0}$, $H_{1}$, or $L_{i}$ in the control law.
+We assume that an estimate of the density matrix $\rho$ is available for feedback.
+In practice, the quantum state is not directly measurable in real time; rather, it must be reconstructed from measurement records using quantum filtering or observer-based techniques, such as those described in [qamar2019observer].
+Remark II.1.
+Measurement and Observer Assumptions.
+The present formulation assumes access to an estimate of density $\rho$ for feedback and neglects the effect of measurement on the system dynamics.
+This corresponds to an ensemble-averaged or mean-field description of the open quantum system and serves as a first step toward evaluating model-free adaptive control at the level of LGKS dynamics.
+In general, a strict separation principle between quantum state estimation and feedback control does not hold for stochastic quantum systems with measurement backaction.
+Incorporating stochastic master equations and measurement-induced backaction into the feedback loop is an important direction for future work.
+$F(\rho,\rho_{\rm{d}})$ Adaptive Controller $\mathcal{L}\mathcal{G}\mathcal{K}\mathcal{S}$ Homodyne Detection State Observer $\rho_{\rm{}_{d}}$ $e$ $u$ $\rho$ $\rho$
+Figure 1: Adaptive feedback control to track desired density matrix $\rho_{\rm d}.$ The estimator, shown in red, consisting of a homodyne detection and a state observer, provides the density matrix $\rho$ for feedback.
 II-A Density Error
-As shown in [jones2024controlling], the Uhlmann–Jozsa fidelity between two density matrices $\rho_{1},\rho_{2}$ is defined as
+As shown in [15], the Uhlmann–Jozsa fidelity between two density matrices $\rho_{1},\rho_{2}$ is defined as
 $F(\rho_{1},\rho_{2})$ $\stackrel{{\scriptstyle\triangle}}{{=}}\left({\rm tr}\,\left(\sqrt{\sqrt{\rho_{1}}\,\rho_{2}\,\sqrt{\rho_{1}}}\right)\right)^{2}.$ (3)
-It follows from Theorem A.5 that $F(\rho_{1},\rho_{2})\in[0,1].$ Furthermore, $F(\rho_{1},\rho_{2})=1$ if and only if $\rho_{1}=\rho_{2}.$ Finally, as shown in [jozsa1994fidelity], in the case where $\rho_{1},\rho_{2}\in\mathbb{C}^{2\times 2},$ the Uhlmann–Jozsa fidelity is given by
+It follows from Proposition A.1 that, if $\rho_{1},\rho_{2}$ are the states of (1), then $F(\rho_{1},\rho_{2})\in[0,1].$ Furthermore, $F(\rho_{1},\rho_{2})=1$ if and only if $\rho_{1}=\rho_{2}.$ Finally, as shown in [16], in the case where $\rho_{1},\rho_{2}\in\mathbb{C}^{2\times 2},$ the Uhlmann–Jozsa fidelity is given by
 $F(\rho_{1},\rho_{2})$ $={\rm tr}\,(\rho_{1}\rho_{2})+2\sqrt{\det(\rho_{1})\det(\rho_{2})}$ (4)
-In this paper, we consider the density error
+In this paper, we define the scalar density-tracking error as
 $e$ $\stackrel{{\scriptstyle\triangle}}{{=}}1-F(\rho,\rho_{\rm{d}})\in[0,1],$ (5)
-where $\rho_{\rm{d}}\in\mathbb{C}^{2\times 2}$ is the desired density
+where $\rho_{\rm{d}}\in\mathbb{C}^{2\times 2}$ is the desired density.
+This formulation of error ensures that the feedback control law is driven by a scalar variable, thereby avoiding any geometric inconsistencies in the simulation of the LGKS system.
+Note that the error $e=0$ when $\rho$ is equal to $\rho_{\rm d}.$
 II-B Adaptive PID Control
-An adaptive proportional-integral-derivative controller can be written as
+In this work, we consider an adaptive proportional-integral-derivative (PID) controller, which can be written as
 $u(t)$ $=k_{\rm p}(t)e(t)+k_{\rm i}(t)\int_{0}^{t}e(\tau){\rm d}\tau+k_{\rm d}(t)\dot{e}(t),$ (6)
-where $k_{\rm p}(t),k_{\rm i}(t),$ and $k_{\rm d}(t)$ are the adaptive proportional, integral, and derivative gains. Note that the controller can be written as
+where $k_{\rm p}(t),k_{\rm i}(t),$ and $k_{\rm d}(t)$ are the adaptive proportional, integral, and derivative gains. Note that the controller can be written in the regressor form as
 $u(t)=\Phi(t)\theta(t),$ (7)
 where
-$\Phi(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}e(t)&\gamma(t)&\dot{e}(t)\end{bmatrix},\quad\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}k_{\rm p}(t)\\ k_{\rm i}(t)\\ k_{\rm d}(t)\end{bmatrix}.$ (8)
-Note that integral state $\gamma(t)$ satisfies $\dot{\gamma}(t)=e(t).$
-The objective of the adaptive controller is to update the controller gains and thus compute the control signal $u(t)$ such that the density error $e(t)$ converges to zero.
-In this paper, we use the continuous-time retrospective cost adaptive control algorithm, described in the next section, to update the controller gains.
+$\Phi(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}e(t)&\gamma(t)&\dot{e}(t)\end{bmatrix},\quad\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}k_{\rm p}(t)\\ k_{\rm i}(t)\\ k_{\rm d}(t)\end{bmatrix}$, (8)
+and the integral state $\gamma(t)$ satisfies $\dot{\gamma}(t)=e(t).$
+The objective of the adaptive controller is to update the gains in $\theta(t)$ online so that the density error $e(t)$ is driven toward zero.
+The gain update law is obtained using the continuous-time retrospective cost adaptive control (RCAC) algorithm described in Section III.
+The PID structure is adopted due to its low-order structure, interpretability, and compatibility with linear parameterizations required by RCAC.
+More general controller parameterizations, such as full-state feedback, can be incorporated within the same adaptive framework when additional measurements or state features are available.
 III Continuous Time Retrospective Cost Adaptive Control
@@
 $y(t)$ $=g(x(t)),$ (10)
-where $x(t)\in\mathbb{C}^{l_{x}}$ is the state, $u(t)\in\mathbb{C}^{l_{u}}$ is the input, $y(t)\in\mathbb{C}^{l_{y}}$ is the measured output, and the vector functions $f$ and $g$ are the dynamics and the output maps.
-Define the performance variable
+where $x(t)$ is the state, $u(t)$ is the input, $y(t)$ is the measured output, and the vector functions $f$ and $g$ are the dynamics and the output maps. Define the performance variable
 $z(t)\stackrel{{\scriptstyle\triangle}}{{=}}y(t)-r(t),$ (11)
 where $r(t)$ is the exogenous reference signal. The objective is to design an adaptive output feedback control law to ensure that $z(t)\to 0.$
+Remark III.1.
+Specialization to LGKS fidelity tracking.
+In the present application, the system state $x(t)$ corresponds to the density matrix $\rho(t)$ governed by the LGKS dynamics (1) with controlled Hamiltonian (2).
+The measured output is chosen as a scalar function of the density matrix, namely the fidelity-based signal $y(t)\stackrel{{\scriptstyle\triangle}}{{=}}F(\rho(t),\rho_{\rm d})$, and the reference is $r(t)\stackrel{{\scriptstyle\triangle}}{{=}}1$.
+Consequently, the performance variable (11) becomes $z(t)=y(t)-r(t)=F(\rho(t),\rho_{\rm d})-1=-e(t)$, where $e(t)$ is the density error defined in (5).
+This choice yields a scalar performance variable that is compatible with the linearly parameterized controller (12).
 Consider a linearly parameterized control law
 $u(t)=\Phi(t)\theta(t),$ (12)
-where the regressor matrix $\Phi(t)\in\mathbb{C}^{l_{u}\times l_{\theta}}$ contains the measured data and the vector $\theta(t)\in\mathbb{C}^{l_{\theta}}$ contains the controller gains to be optimized.
+where the regressor matrix $\Phi(t)$ contains the measured data and the vector $\theta(t)$ contains the controller gains to be optimized.
 Various linear parameterizations of MIMO controllers are described in [goel_2020_sparse_para].
@@
 where $\hat{\theta}(t)$ is the controller gain to be optimized and the filtered regressor $\Phi_{\rm f}(t)$ and the filtered control $u_{\rm f}(t)$ are defined as
-$\Phi_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[\Phi(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (14)
-$u_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[u(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (15)
+$\Phi_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[\Phi(t)\right],$ (14)
+$u_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[u(t)\right],$ (15)
 where $G_{\rm f}(s)$ is a dynamic filter.
@@
 $u_{\rm f}$ $=C_{\rm f}x_{u}+D_{\rm f}u.$ (19)
+Remark III.2.
+Filter states and notation.
+The states $x_{\phi}$ and $x_{u}$ appearing in the filter realizations below are internal states of the dynamic filter $G_{\rm f}(s)$ driven by $\Phi(t)$ and $u(t)$, respectively.
+These filter states are auxiliary variables used only to compute $\Phi_{\rm f}(t)$ and $u_{\rm f}(t)$ and are not related to the LGKS state $\rho(t)$.
 Next, define the retrospective cost
-$J(t,\hat{\theta})$ $=\int_{0}^{t}\Big{(}e^{-\lambda(t-\tau)}\hat{z}(\tau)^{\rm T}R_{z}\hat{z}(\tau)$
-$\quad+(\Phi(\tau)\hat{\theta})^{\rm T}R_{u}(\Phi(\tau)\hat{\theta})\Big{)}{\rm d}\tau$
-$\quad+e^{-\lambda t}\hat{\theta}^{T}R_{\theta}\hat{\theta},$ (20)
-where $R_{z},R_{u},$ and $R_{\theta}$ are positive definite weighting matrices of appropriate dimensions, and $\lambda<0$ is the exponential forgetting factor.
+[…]
+$\quad+(\Phi(\tau)\hat{\theta})^{\rm T}R_{u}(\Phi(\tau)\hat{\theta})\Big){\rm d}\tau+e^{-\lambda t}\hat{\theta}^{T}R_{\theta}\hat{\theta}$, (20)
+where $R_{z},R_{u},$ and $R_{\theta}$ are positive definite weighting matrices of appropriate dimensions, and $\lambda>0$ is the exponential forgetting factor.
+The retrospective cost (20) defines an online performance objective for updating the parameter vector $\theta(t)$ based on measured data.
+We emphasize that minimizing (20) does not, in general, imply closed-loop stability or guarantee that $z(t)\to 0$ for nonlinear systems.
+In particular, for the LGKS dynamics considered in this paper, the tracking behavior reported in Section IV is empirical and depends on the system, the controller parameterization, and the choice of hyperparameters.
 Proposition III.1.
-Consider the cost function $J(t,\hat{\theta})$ given by (20). For all $t\geq 0,$ define the minimizer of $J(t,\hat{\theta})$ by
-$\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\underset{\hat{\theta}\in{\mathbb{R}}^{l_{\theta}}}{\operatorname{argmin}}J(t,\hat{\theta}).$ (21)
+[…]
+[…]
 Then, for all $t\geq 0,$ the minimizer satisfies
@@
 Note that the cost function (20) can be written as
-$J(t,\hat{\theta})$ $=\hat{\theta}^{T}A(t)\hat{\theta}+2\hat{\theta}^{T}b(t)+c(t),$ (24)
+[…]
 where
@@
 where $A(0)=R_{\theta}$ and $b(0)=0.$
-Next, define, for all $\geq 0,$ $P(t)\stackrel{{\scriptstyle\triangle}}{{=}}A(t)^{-1}.$ Using the fact that $P(t)A(t)=I,$ it follows that
+Next, define, for all $t\geq 0,$ $P(t)\stackrel{{\scriptstyle\triangle}}{{=}}A(t)^{-1}.$ Using the fact that $P(t)A(t)=I,$ it follows that
 $\dot{P}(t)=-P(t)\dot{A}(t)P(t),$
@@
 where $\theta(t)$ is given by (22).
+For the adaptive PID structure in Section II, the regressor $\Phi(t)$ and parameter vector $\theta(t)$ in (12) correspond to the fidelity-based features and gains defined in (6).
+Thus, the update laws (22) provide an online method to adjust $(k_{\rm p}(t),k_{\rm i}(t),k_{\rm d}(t))$ using only the measured fidelity-based error signal and its integral/derivative features, without requiring explicit knowledge of $H_{0}$, $H_{1}$, or $L_{i}$ in the control law.
 IV Numerical Simulation
-In this paper, we consider a two-level system, with dynamics given by the LGKS (1), with a single damped term, that is, $n=1$. The control system architecture is shown in Figure 1.
-$F(\rho,\rho_{\rm{d}})$$\begin{array}[]{c}\text{Adaptive}\\ \text{Controller}\end{array}$ $\mathcal{L}\mathcal{G}\mathcal{K}\mathcal{S}$$\rho_{\rm{d}}$$e$$u$$\rho$
-Figure 1: Adaptive control architecture to track desired density matrices in the LGKS system.
-To simulate the LKGS equation, we set
+In this paper, we consider a two-level LGKS system (1) with a single dissipative channel, that is, $m=1$.
+The control system architecture is shown in Fig.
+1.
+Note that, for conceptual clarity, the observer is included in the block diagram; however, in the numerical studies, we assume access to an estimate of the density matrix $\rho$ for feedback computation.
+In practice, the quantum state is not directly measurable in real time and must be reconstructed from measurement records using quantum filtering or observer-based techniques; a representative observer design is described in [qamar2019observer].
+In this preliminary study, the observer dynamics and measurement backaction are not explicitly modeled, and the focus is on evaluating the closed-loop behavior of the adaptive controller under idealized assumptions about state access.
+To simulate the LGKS equation, we set
 $H_{0}=0.5\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},H_{1}=0.5\begin{bmatrix}0&1\\ 1&0\end{bmatrix},L_{1}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$, (26)
 where $H_{0}$ and $H_{1}$ are Pauli matrices that, as described in [d2001optimal, kamakari2022digital], satisfy the controllability conditions in the problem of controlling a two-level quantum system, and $L_{1}$ is selected as a non-Hermitian matrix, as suggested in [verstraete2009quantum, schlimgen2022quantum].
-The initial density matrix $\rho(0)$ is assumed to be
-$\rho_{0}$ $=\begin{bmatrix}0.4&0.1+0.3\imath\\ 0.1-0.3\imath&0.6\end{bmatrix}.$ (27)
-Note that $\rho(0)$ is randomly generated such that ${\rm tr}\,\rho(0)=1$ and it is Hermitian.
-As shown in [moses2015creation] and [xu2021multicriticality], we consider the problem of moving the quantum state $\rho(0)$ to a low and a high desired entropy states $\rho_{\rm d},$ where the Von Neumann entropy of a quantum state $\rho$ is given by
+The initial density matrix is chosen as
+$\rho_{0}$ $=\begin{bmatrix}0.4&0.1+0.3\imath\\ 0.1-0.3\imath&0.6\end{bmatrix}$, (27)
+which is a randomly generated Hermitian, positive semidefinite matrix with unit trace.
+Following [22, 34], we consider both low-entropy and high-entropy target states $\rho_{\rm d}$, where the von Neumann entropy of a quantum state $\rho$ is defined as
 $S(\rho)\stackrel{{\scriptstyle\triangle}}{{=}}-{\rm tr}\,(\rho\,\rm{ln}(\rho)).$ (28)
-First, we consider the problem of driving the quantum state $\rho(t)$ to
+Achieving or maintaining low-entropy states is often a central objective in quantum information processing, since low entropy corresponds to high purity and coherence, which are essential resources for quantum computation, communication, and sensing.
+Because environmental interactions and decoherence naturally increase entropy, feedback control is required to counteract this drift and preserve useful quantum states.
+Conversely, the ability to track high-entropy target states is also relevant in many physical scenarios, since open quantum systems are often driven toward mixed or thermal equilibrium states by their environment.
+IV-A Hyperparameter Tuning
+First, we specify the low-entropy target state
 $\rho_{\rm d}=\begin{bmatrix}0.8571&0.2857+0.1429\imath\\ 0.2857-0.1429\imath&0.1429\end{bmatrix}$. (29)
-Note that $S(\rho_{\rm d})=0.1013,$ and thus, this desired state is a low-entropy state.
-Furthermore, note that $\rho_{\rm d},$ given by (29) is an equilibrium point of (1) in the case where $u(t)\equiv 1.$
-With the desired density $\rho_{\rm d}$ given by (29) as the reference for the controller, we first tune the RCAC hyperparameters.
-In particular, in RCAC, we set $R_{z}=1,$ $R_{u}=1,$ and $\lambda=0.01$; and optimize $P_{0}$ and the filter coefficient $\beta,$ where $G_{\rm f}(s)=\dfrac{1}{s+\beta},$ as described below.
-We optimize only $P_{0}$ and the filter coefficient since it has been observed in several numerical experiments that the choice of $P_{0}$ and the filter coefficient has the most dominant impact on the closed-loop performance.
-For each pair of $(P_{0},\beta)\in\{(10^{-5},10^{-4},\ldots,10^{10})\times(0,1,2,5,100,2000)\},$ we apply the RCAC algorithm to update the adaptive PID controller to drive the quantum state from $\rho_{0}$, given by (27) to $\rho_{\rm d}$, given by (29).
+Note that $S(\rho_{\rm d})=0.1013,$ and thus, this desired state is a low-entropy state. Furthermore, $\rho_{\rm d},$ given by (29), is an equilibrium point of (1) in the case where $u(t)\equiv 1.$
+With the desired density $\rho_{\rm d}$ given by (29) as the reference for the controller, we tune the RCAC hyperparameters.
+In particular, we set $R_{z}=1,$ $R_{u}=1,$ and $\lambda=0.01,$ and optimize $P_{0}$ and the filter coefficient $\beta,$ where $G_{\rm f}(s)=\dfrac{1}{s+\beta}.$ We restrict the optimization to $P_{0}$ and $\beta,$ as numerical experiments consistently indicate that these parameters are the most dominant factors in the closed-loop performance.
+For each pair of $(P_{0},\beta)\in\{(10^{-5},10^{-4},\ldots,10^{10})\times(0,1,2,5,100,2000)\},$ the RCAC algorithm is applied to update the adaptive PID controller to drive the quantum state from $\rho_{0}$, given by (27), to $\rho_{\rm d}$, given by (29).
 Each simulation is run for 200 seconds.
@@
 is computed.
-Figure 2 shows the contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The tuning pair with the minimum value of $J_{h}$ is selected for all the numerical simulations in the rest of the paper.
-In particular, in all the examples, we set $P_{0}=10^{-3}\times I_{3},\beta=2000.$
-Figure 2: Contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The pair $(P0,\beta)$ that minimizes $J_{h}$ is used in all numerical simulations.
-IV-A Low-entropy Density
-With the selected RCAC hyperparameters, Figure 3 shows the absolute value of the density error $e(t)$ between the system state and the desired low-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller.
-Figure 4 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state.
-Figure 5 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
-Note that the quantum state $\rho$ satisfies $x=2\,\rm{Real}(\rho_{1,2}),$ $y=2\,\rm{Imag}(\rho_{1,2}),$ and $z=\rho_{1,1}-\rho_{2,2}$ on the Bloch sphere [wie2020two].
+Figure 2 shows the contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The tuning pair with the minimum value of $J_{h}$ is selected for all numerical simulations presented in the remainder of the paper.
+In particular, in all the examples, we set $P_{0}=10^{-3}\times I_{3}$ and $\beta=2000.$ We emphasize that the hyperparameter tuning procedure outlined above is coarse and heuristic in nature, and is intended to identify a representative near-optimal parameter set for the examples considered here.
+Figure 2: Contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The pair $(P_{0},\beta)$ that minimizes $J_{h}$ is used in all numerical simulations.
+IV-B Low-Entropy Density Tracking
+With the selected RCAC hyperparameters, Fig.
+3 shows the absolute value of the density error $e(t)$ between the detected system state and the desired low-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ generated by the adaptive controller.
+Fig.
+4 shows the real and imaginary parts of the quantum state $\rho(t)$ and the von Neumann entropy of the quantum state.
+Fig.
+5 shows the trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
+Note that the quantum state $\rho$ satisfies $x=2\,\rm{Real}(\rho_{1,2}),$ $y=2\,\rm{Imag}(\rho_{1,2}),$ and $z=\rho_{1,1}-\rho_{2,2}$ on the Bloch sphere [33].
 Figure 3: Low-entropy density tracking.
-The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
+The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ generated by the adaptive controller.
 Figure 4: Low-entropy density tracking.
-The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color.
+The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ The desired state components are shown with dashed lines, while the system response is shown with solid lines of the corresponding color.
 The third subplot shows the entropy of the quantum state.
-Figure 5: Low-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
-IV-B High-entropy Density
-Next, we consider the problem of driving the quantum state $\rho(t)$ to
+Figure 5: Low-entropy density tracking.
+Trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
+Note that the trajectory remains inside the Bloch ball, consistent with mixed-state evolution.
+IV-C High-Entropy Density Tracking
+Next, we consider the problem of driving the quantum state $\rho(t)$ to the high-entropy target
 $\rho_{\rm d}=\begin{bmatrix}0.5168&0.0971+0.0460i\\ 0.0971-0.0460i&0.4832\end{bmatrix}$. (31)
-Note that $S(\rho_{\rm d})=0.6693,$ and thus, this desired state is a high-entropy state.
-Furthermore, note that $\rho_{\rm d},$ given by (31) is an equilibrium point of (1) in the case where $u(t)\equiv 10.$
-Figure 6 shows the absolute value of the density error $e(t)$ between the system state and the desired high-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller.
-Figure 7 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state.
-Figure 8 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
-The two examples described above suggest that the adaptive PID controller updated by the RCAC algorithm can effectively track an arbitrary desired density matrix $\rho_{\rm d}$ without requiring any model of the underlying dynamics of the density matrix.
+Note that $S(\rho_{\rm d})=0.6693,$ and thus, this desired state is a high-entropy state. Furthermore, $\rho_{\rm d},$ given by (31), is an equilibrium point of (1) in the case where $u(t)\equiv 10.$
+Figure 6 shows the absolute value of the density error $e(t)$ between the detected system state and the desired high-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control input $u$ generated by the adaptive controller.
+Figure 7 shows the real and imaginary parts of the quantum state $\rho(t)$ and the von Neumann entropy of the quantum state.
+Figure 8 shows the trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
 Figure 6: High-entropy density tracking.
-The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
+The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ generated by the adaptive controller.
 Figure 7: High-entropy density tracking.
-The first two subplot shows the real and imaginary part of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color.
+The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ The desired state components are shown with dashed lines, while the system response is shown with solid lines of the corresponding color.
 The third subplot shows the entropy of the quantum state.
-Figure 8: High-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
+Figure 8: High-entropy density tracking.
+Trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
+Note that the trajectory remains inside the Bloch ball, consistent with mixed-state evolution.
+IV-D Shot Noise
+Finally, we revisit the low-entropy target $\rho_{\rm d}$ given by (29) and examine the effect of measurement noise on the feedback loop.
+Shot noise in the feedback detection is modeled as additive white noise, following [niebauer1991nonstationary].
+Specifically, at each discrete sampling instant, a perturbation $N\in\mathbb{C}^{2\times 2}$ is drawn with entries distributed as $\mathcal{N}(0,\sigma^{2})$, and the noisy state is constructed as
+$\rho_{\text{noisy}}=\frac{\rho+\tfrac{1}{2}(N+N^{\dagger})}{{\rm tr}\,\!\left(\rho+\tfrac{1}{2}(N+N^{\dagger})\right)}.$ (32)
+Hermiticity is enforced by symmetrization, and physical validity is preserved by trace normalization, ensuring that $\rho_{\text{noisy}}$ remains a valid density matrix under white noise of strength $\sigma=5\times 10^{-3}$.
+Remark IV.1.
+Noise model and feedback-only corruption.
+The noisy state $\rho_{\text{noisy}}$ is used only in the feedback path for computing the fidelity-based error signal.
+The underlying system evolution is still governed by the deterministic LGKS dynamics driven by the true state $\rho(t)$.
+Thus, the injected noise models the corruption of the measurement or state estimate rather than physical noise acting directly on the quantum dynamics.
+Figure 9 shows the absolute value of the density error $e(t)$ between the detected system state with shot noise and the desired state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ generated by the adaptive controller.
+Figure 10 shows the real and imaginary parts of the quantum state $\rho(t)$ and the von Neumann entropy of the quantum state.
+Figure 11 shows the trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
+The three examples described above suggest that the adaptive PID controller updated by the RCAC algorithm can effectively track a desired density matrix $\rho_{\rm d}$ in the presence of moderate measurement noise, without requiring a model of the underlying LGKS dynamics in the control law.
+Figure 9: Shot noise case tracking.
+The first subplot shows the absolute value of the density error $e(t)$ between the detected system state with shot noise and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ generated by the adaptive controller.
+Figure 10: Shot noise case tracking.
+The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ The desired state components are shown with dashed lines, while the system response is shown with solid lines of the corresponding color.
+The third subplot shows the entropy of the quantum state.
+Figure 11: Shot noise case tracking.
+Trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
+Note that the trajectory remains inside the Bloch ball, consistent with mixed-state evolution.
 V Conclusions
-This paper presented an online learning-based, model-free adaptive controller for tracking a desired density matrix in a two-level Lindblad-Gorini-Kossakowski-Sudarshan quantum system.
-In particular, the adaptive controller uses the retrospective cost adaptive control algorithm.
-Since the classical RCAC was developed for a discrete-time system, the paper presented a continuous-time extension of RCAC.
-An adaptive PID controller structure was chosen to generate a control signal that preserves the LKGS solution’s geometric structure.
-Uhlmann-Jozsa fidelity was used as an error metric to drive the adaptive PID controller.
-Numerical simulations show that the proposed controller effectively tracks desired low and high-entropy states without requiring a model of the underlying system dynamics.
-Future work will extend the technique to the case where the input is a Hermitian matrix and the density coherence is used as an error metric.
-This will allow higher-dimensional parameterization of the adaptive PID controller, which may yield better tracking performance.
-[…]
+This paper presented an online learning-based adaptive control framework for tracking a desired density matrix in a two-level Lindblad–Gorini–Kossakowski–Sudarshan quantum system, in which the feedback control law does not require prior knowledge of the system Hamiltonian or dissipative operators.
+In particular, the adaptive controller is based on the retrospective cost adaptive control (RCAC) algorithm, and a continuous-time formulation is employed to align with the LGKS dynamics.
+An adaptive PID controller structure is used to generate the control signal while preserving the geometric structure of the LGKS evolution.
+Uhlmann–Jozsa fidelity is adopted as a scalar performance metric to drive the adaptation.
+Numerical simulations demonstrate effective tracking of both low-entropy and high-entropy target states under idealized state-access assumptions and in the presence of moderate measurement noise.
+Future work will extend this framework to matrix-valued control inputs and alternative quantum error metrics, such as coherence-based measures, enabling richer controller parameterizations and evaluation on higher-dimensional quantum systems.
+Additional directions include incorporating stochastic measurement backaction, discrete-time feedback implementations, and theoretical analysis of closed-loop performance guarantees.
 Appendix
@@
 Definition A.1.
-Let $A\in{\mathbb{C}}^{n\times n}.$ $A$ is a Hermitian matrix if $A=\overline{A^{\rm T}},$ where $\overline{A^{\rm T}}$ is the conjugate transpose of $A.$
+Let $\in{\mathbb{C}}^{n\times n}. AA$ is a Hermitian matrix if $A=\overline{A^{\rm T}},$ where $\overline{A^{\rm T}}$ is the conjugate transpose of $A.$
 The conjguate transpose of $A$ is denoted by $A^{\rm H}.$ If $A$ is Hermitian, then $A=A^{\rm H}.$
@@
 Fact A.1.
-Let $A,B\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $-\imath[A,B]$ is Hermitian.
+Let $A,B\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $\imath[A,B]$ is Hermitian.
 Proof.
 Note that $(-i[A,B])^{\rm H}=(-\imath)^{\rm H}[A,B]^{\rm H}=\imath\left(AB-BA\right)^{\rm H}=\imath(A^{\rm H}B^{\rm H}-B^{\rm H}A^{\rm H})=\imath(BA-AB)=-\imath(AB-BA)=-\imath[A,B]$.
-Thus, $-\imath[A,B]$ is Hermitian.
+Thus, $\imath[A,B]$ is Hermitian.
 ∎
@@
 which completes the proof. ∎
-Fact A.5.
+Proposition A.1.
 Let $\rho\in\mathbb{C}^{n\times n}$ and $\sigma\in\mathbb{C}^{n\times n}$ be two Hermitian PSD matrices with traces equal to $1$. Then,
-$1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$
+$1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0.$
 Proof.
 Since $\rho$ and $\sigma$ are PSD Hermitian.
-Then, $\rho^{{\rm H}}=\rho$, $\sigma^{{\rm H}}=\sigma$, $x^{{\rm H}}\rho x\geq 0$ and $x^{{\rm H}}\sigma x\geq 0$ for all $x\in\mathbb{C}.$ Furthermore, the unique square root of a PSD Hermitian matrix is also Hermitian and PSD [horn2012matrix, p 440].
+Then, $\rho^{{\rm H}}=\rho$, $\sigma^{{\rm H}}=\sigma$, $x^{{\rm H}}\rho x\geq 0$ and $x^{{\rm H}}\sigma x\geq 0$ for all $x\in\mathbb{C}.$ Furthermore, the unique square root of a PSD Hermitian matrix is also Hermitian and PSD [13, p 440].
 This implies that, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\left(\sqrt{(}\rho)\sigma\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\sigma^{{\rm H}}\sqrt{\rho}^{{\rm H}}=\sqrt{\rho}\sigma\sqrt{\rho}.$ Define, $y\stackrel{{\scriptstyle\triangle}}{{=}}\sqrt{\rho}x\in\mathbb{C}^{n}.$ Then, $x^{{\rm H}}\sqrt{\rho}\sigma\sqrt{\rho}x=y^{{\rm H}}\sigma y\geq 0,$ which implies that $\sqrt{\rho}\sigma\sqrt{\rho}\geq 0.$ Therefore, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)$ is Hermitian, and $\sqrt{\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)}$ is Hermitian and unique.
 From the Cauchy-Schwarz inequality and the cyclic property for traces, $|{\rm tr}\,(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}})|^{2}=|{\rm tr}\,(\sqrt{\sqrt{\rho}\sqrt{\rho}\sigma})|^{2}=|{\rm tr}\,(\sqrt{\rho\sigma})|^{2}\leq{\rm tr}\,(\rho){\rm tr}\,(\sigma)\leq 1$.
-Thus, $1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\rho_{\rm{d}}\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$
+Thus, $1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\rho_{\rm{d}}\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0.$